OpenStudy (rayford):

x^2-7x+6 use the descrimitave formula to find how many and what types of roots ill have

OpenStudy (anonymous):

Do you know the discriminant formula?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

The discriminant formula is \[\sqrt{b^2-4ac}\] you might notice that this is from the quadratic formula \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

first use your knowledge of a quadratic equation. It can have 2 possible zeroes

OpenStudy (rayford):

nevermind thank you!

OpenStudy (anonymous):

alright i'm guessing you just didn't understand the question but understood how to do the problem?